OpenStudy (anonymous):

Find the sum of the series shown below. 3 + (-6) + (-15) + ... + (-348)

OpenStudy (anonymous):

no terms missing, like 9? or 12

OpenStudy (anonymous):

Thats if its geometric.

OpenStudy (anonymous):

right,,,, o if you add all them together ( even though they are negative ) what is the sum.

OpenStudy (anonymous):

what do you mean right...

OpenStudy (anonymous):

So is there supposed to be a 9 and 12?

OpenStudy (anonymous):

add all the terms from , 3,-6,-15,-24,-33,-42,-51,-60 all the way to -348... whats the sum of that series

OpenStudy (cwrw238):

could it be arithmetic with common difference -9?

OpenStudy (anonymous):

and right means im agreeing with you

OpenStudy (anonymous):

Good work cwrw238. I have to go back to class now.

OpenStudy (anonymous):

sum = (3-348)39/2

OpenStudy (cwrw238):

sum n terms = n/2 [ a + l] but how do we find n?

OpenStudy (anonymous):

n = 39

OpenStudy (anonymous):

ok since n=39 the equation should be 39{{3+(-348)} ------------ 2

OpenStudy (cwrw238):

348 - 6 = 342 342/9 = 38 + 1 = 39 - oh yes sum = (39/2) (3 - 348) but that s not a whole number

OpenStudy (cwrw238):

so in n = 39 correct?

OpenStudy (cwrw238):

i think its 40 because if we take 3, - 6, - 15 15-6 / 9 = 1 so we need to add 2

OpenStudy (cwrw238):

so sum = 20 ( 3 - 348) = 20 * -345 = -6900

OpenStudy (cwrw238):

yep - thats correct devin

OpenStudy (cwrw238):

i'm convinced of that because if we take say 4 terms 3 -6 -15 - 24 the sum of this is -42 use formula sum = (4/2)*(3 - 24) = 2 * -21 = -42 7 terms : 3,-6,-15,-24,-33,-42,-51 the sum is -168 use formula (7/2)*(3 - 51) = (-48*7 ) / 2 = -168